[Ncep.nhc.nco_contacts] TAFB Shift Log -- Technical Edition -- Thu Jul 28, 2016

TAFB Shift Log nhc.tsbadmin at noaa.gov
Fri Jul 29 01:05:34 UTC 2016


TAFB Shift Log for Thu Jul 28 2016 21:05:31 GMT-0400 (EDT)

Shift: Swing Shift

Pacific Forecaster: H Cobb

Atlantic Forecaster: A Levine

Surface Analyst: E Rivera-Acevedo

Technical Issues Summary:

GFE/NCP issues:

N-AWIPS issues:

Other issues: GOES-W imagery stopped ingesting at 2330Z. Called NCEP SDM at  
0025Z to inquire. Severe thunderstorms in the Wallops lsland area caused  
imagery dropouts beginning at 2312Z. Additionally there were hung processes  
on the NCEP satellite server. The SDM cleared the processes at 0050Z. As of  
0100Z still awaiting arrival of new imagery.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160729/4dca2713/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list